find two consecutive positive even numbers such that the sum of their squares is 164.
Answer by Alan3354(69443)

8 and 10
------------
Do I need to work it out?
x^2 + (x+2)^2 = 164
2x^2 + 4x + 4 = 164
x^2 + 2x - 80 = 0
(x-8)*(x+10) = 0
x = 8
x+2 = 10